### v3

# DAG-Plan: Generating Directed Acyclic Dependency Graphs for Dual-Arm Cooperative Planning

Abstract
Dual-arm robots offer enhanced versatility and efficiency over single-arm counterparts by enabling concurrent manipulation of multiple objects or cooperative execution of tasks using both arms. However, the coordination of dual-arm systems for long-horizon tasks continues to pose significant challenges, stemming from the intricate temporal and spatial dependencies among sub-tasks, necessitating intelligent decisions regarding the allocation of actions between arms and their optimal execution order. Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations to use large language models (LLMs) generate task sequence with linear temporal dependency, failing to fully leverage the capabilities of dual-arm systems. To address this limitation, we introduce DAG-Plan, a structured task planning framework tailored for dual-arm robots. DAG-Plan harnesses LLMs to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG). Critically, DAG-Plan dynamically assigns these sub-tasks to the appropriate arm based on real-time environmental observations, enabling parallel and adaptive execution. We evaluate DAG-Plan on the Dual-Arm Kitchen Benchmark, comprising 5 sequential tasks with 44 sub-tasks. Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate linear task sequence, achieving 52.8% higher efficiency compared to the single-arm task planning and 48% higher success rate of the dual-arm task planning. Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time. More demos and information are available on https://sites.google.com/view/dag-plan.
Index Terms:
Dual-arm Robots, Task Planning, LLMs
I Introduction
Achieving effective bimanual coordination in robotics is challenging due to the complexities and long-horizon of dual-arm operations, requiring precise spatial and temporal coordination [1, 2]. While humans effortlessly coordinate their hands in daily long-horizon tasks, replicating such coordination in robots presents significant challenges. Traditional method employed hand-designed primitives to manage the movements of dual robotic arms [3, 4]. Whereas these methods often fall short in long-horizon task as they lack the flexibility required for adaptive task execution. Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [5, 6]. By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to long-horizon scenarios and tasks in the wild [7, 8].
Existing LLMs planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills. While some studies [9, 10] have employed dual-arm robots as test platforms, these still engage only one arm at a time or rely on predefined bimanual tasks, leading to inefficiencies. Several multi-agent planning methods  [11, 12, 13] have been proposed, which assign tasks to each robot through independent task decomposition or interactions among multiple robots, significantly enhancing performance in multi-robot collaborative tasks. These methods share a common characteristic: they all utilize LLMs to generate task sequences with linear temporal dependencies.
Generating dual-arm planning schemes as task sequences with LLMs faces significant challenges, primarily three gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex. Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed in non-iterative method, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution; 3) The iterative approach, while capable of interacting with the environment and achieving a higher success rate compared to non-iterative methods, necessitates querying the LLM for each decision, resulting in significant token consumption and time costs.
To address these challenges, we introduce DAG-Plan, a structured task planning framework that leverages the capabilities of LLMs. We leverage Directed Acyclic Graph (DAG) as a task graph for dual-arm task planning. A DAG represents each complex task as actionable sub-tasks, with nodes indicating these sub-tasks and directed edges defining explicit temporal dependencies. Firstly, we utilize LLMs to generate the DAG, decomposing complex tasks into nodes, each associated with a specific type and the number of arms required for execution. Subsequently, the DAG enters the task planning inference process. DAG-Plan uses this temporal dependency information and node types to determine priority candidate nodes and common candidate nodes, assigning them to the left and right arms. DAG-Plan checks the feasibility and calculates the cost of combinations of left and right arm candidate nodes based on the environmental state, adaptively executing sub-tasks that are easier and closer to perform.
Our main contributions are summarized as follows:
  • We identify the limitations of planning methods which generate linear task sequences when applied to dual-arm robots, particularly their inefficiency in handling complex temporal dependencies and their inability to adaptively interact with the environment during task execution.
  • We present DAG-Plan, an efficient cooperative task planning framework for dual-arm robots. This framework represents decomposed sub-tasks as a DAG with LLMs and dynamically assigns sub-tasks to the appropriate arm based on the real-time environment state.
  • Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods. DAG-Plan achieves 52.8% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan. Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates 48% higher success rate. Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time.
II Related Works
II-A Task Planning with LLMs
LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language. Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [14, 9, 10, 15, 11, 12, 13]. However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep. Additionally, these studies [9, 11] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability. Multi-agents planning is the closest to dual-arm planning. Twostep [11] decomposes task into two independent tasks. RoCo [12] and DABICO [13] assign each robot with a LLM to discuss and collectively reason task strategies. These methods either fail to consider the constraints of dual-arm robot or are limited by the linear dependencies of task sequence.
II-B Dual-arm Robot Manipulation
Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [16, 17] with fixed process operations and domestic settings [18, 19] with single-skill operations. With the development of methods such as motion planning [17, 19], foundation model [20, 21], reinforcement learning [22, 23], and imitation learning [24, 25], dual-arm robots can perform many human-like operations at the skill level. However, they still lack the ability to autonomously execution in long-horizon task. The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously execution.
II-C Structured Task Decomposition (STD)
STD involves breaking down a complex task into a DAG. Previous methods for STD, such as Crowd-Sourced STD [26, 27] and Query-based STD [28, 29], were limited by data availability. However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD. TaskLAMA [30] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies. In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
III Methods
We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node. The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed. This ensures that dependencies are meticulously adhered to, allowing for efficient task execution. The robot dynamically assesses the state of its environment and the status of its arms to select next optimal tasks from the graph. This ongoing selection process prioritizes activities, keeping both arms continuously engaged. An overview of DAG-Plan pipeline is illustrated in Figure 1.
Figure 1: An overview of DAG-Plan. The DAG-Plan generates a DAG based on human instruction and environmental description. It checks the graph’s completeness and reflects the LLM to regenerate if incomplete. Once a valid DAG is obtained, DAG-Plan performs task inference to identify executable candidate nodes. The occupied arm and free arm are assigned priority candidate nodes and common candidate nodes respectively. The framework then evaluates all candidate combinations for feasibility and cost. DAG-Plan selects the nodes with the lowest cost and employs skill in library for execution. DAG-Plan updates the graph, iterating inference until the DAG is fully executed.
III-A Directed Acyclic Sub-task Dependency Graph Generation
In prior research, task planning for robots typically involves generating a linear sequence of sub-tasks. However, this model falls short in scenarios involving dual-arm robots, where the capability for parallel task execution can significantly enhance operational efficiency. By coordinating both arms, many tasks can be performed concurrently, which the linear model does not exploit fully.
To address this limitation, we propose a novel approach using LLMs to generate optimized dual-arm plans. Our method involves decomposing complex tasks into a dual-arm task graph instead of a linear sequence. This graph better represents the complex temporal dependencies of bimanual operations. We define the graph as $G=(V,E,T,N)$, where $V$ denotes the tasks, $E$ represents the dependencies, $T$ categorizes the task types, and $N$ specifies the arm requirements. Each vertex $v_{i}\in V$ corresponds to a specific sub-task, and each directed edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that $v_{i}$ must be completed before $v_{j}$. The task types include: 1) Occupy, where tasks involve the engagement of the robot’s gripper and the arm will be occupied after execution, typically for grasping or holding an object; 2) Tool use, which refers to tasks that require the use of a tool, remaining in the gripper throughout the operation; 3) Release, for tasks where an object is released from the gripper, often associated with placement or release into a specific location; 4) Operate, denoting general operational tasks that leave the gripper free post-completion, where the arm will be occupied during execution and released afterward; and 5) Complete, which marks the end of all tasks, represented as the terminal node in the graph. This classification aids in specifying the nature of the task and the number of arms required, thereby enabling a more sophisticated and efficient planning strategy tailored for dual-arm robotic systems. The arm number of node $n_{i}\in N$ represents the arm number of node needed. The occupy-release pairs are crucial structures in dual-arm task graphs. An occupy-release pair mainly consists of an occupy node as start point and a release node as end point, with potentially several tool use nodes in between. A complete occupy-release pair ensures that the robot arm is not continuously occupied and prevents placing an object without first grasping it. Additionally, the dual-arm task graph should be a fully connected DAG. After generating the dual-arm task graph, we check it for completeness. If the graph contains incomplete grasp-release pairs or is not fully connected, we reflect the LLMs to regenerate the task graph.
III-B Task Planning Inference with Generated Directed Acyclic Graph
After generating the dual-arm task graph, the planning process enters the inference phase. This phase utilizes the task graph and the observed state to dynamically refine the planning of robotic arm operations. It selects and executes sub-tasks that are executable and have the lowest cost, based on the task graph and the current environment state. As shown in Figure 2, we provide a detailed and specific illustration of the task planning inference process. We first identify the two types of candidate nodes in DAG-Plan: common candidate nodes and priority candidate nodes. The common candidate nodes include those that can be executed when the robotic arm is in the free state. The priority candidate nodes include subsequent nodes when the arm is already engaged in an occupy-release pair. The common candidate nodes are initially selected by identifying nodes within graph that no other nodes point to.
During execution, once a node completes its operation, it and its associated edges are removed from the graph. This unlocks nodes that are dependent solely on the executed node. If the node executed involves operations like operate or release, the corresponding arm becomes free, and any dependent nodes unlocked by this action are added to the common candidates. Conversely, if the executed node involves actions like occupy or tool use, where the arm remains occupied, any dependent nodes in this occupy-release pair are unlocked and placed into the priority candidates for that arm. When the priority candidate nodes for a specific arm are not empty, the arm must select a sub-task from these priority candidate nodes. This ensures that the arm’s next tasks are aimed at completing actions necessary to free up the arm. When there are no more priority candidates for an arm, that arm is considered free and can select tasks from the common candidates. This strategic selection and execution framework ensures efficient operation and task handling by the dual-arm robotic system.
Figure 2: The process of Task Planning Inference. In task 2 “clean the table (Hard)”, DAG-Plan initializes common candidate nodes based on the DAG. It evaluates node combinations, checks feasibility, and calculates costs. The right arm is selected to grasp apple and the left to open drawer. After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for right arm. In stage 2, right arm is assigned corresponding priority candidate nodes, checked and left arm still selected node in common candidate nodes with empty priority candidate nodes. The task graph and nodes are updated again. In stage 3, left arm put mug into drawer and right arm grasp lemon.
Once we obtain the candidate nodes for each robotic arm, we generate all possible combinations of left and right arm candidate nodes. These combinations are then checked for feasibility, and any pairs that fail the checks are removed from the candidate set. There are three checks in total. The first check involves verifying the presence of an occupy node within the candidates. If an occupy node is found, we further examine whether its successor nodes contain dependencies that require other conditions to be met first. This could lead to prolonged arm occupancy even dead-lock, thereby decreasing operational efficiency. The second check evaluates the distance between the target objects assigned to the left and right arms. If the distance exceeds a predefined threshold, simultaneous operation of both arms becomes infeasible. The third check analyzes the relative positions of the target locations for each arm. If the target position for the left arm is located to the right of the right arm’s target or is too far away, the arms may collide or fail to reach their targets. To acquire the object positions necessary for decision-making, DAG-Plan utilizes Grounding DINO[31] and SAM2[32] to obtain the point clouds corresponding to the specified objects, thereby determining their locations. For specific categories of objects such as bottles and bowls, the category-specific 6D pose estimation model SAR-Net[33] is employed to estimate their poses.
After completing the checks, we calculate the cost of the left and right hand candidate nodes based on the environment state. We aim for the target objects to be close to the robotic hands facilitating dual-arm operations. The pair with the lowest combined cost is selected for execution. The cost $J$ is represented as:
$\footnotesize J=\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}}).$
III-C Sub-tasks Execution with Foundation Model
Once the sub-tasks to be executed are determined, the robot needs to perform the corresponding actions to bridge the gap between textual instructions and the physical environment. The collective perceptual inferences specific to objects, along with physical insights and parameters for manipulation, are methodically organized and converted into a structured format of executable action code.
III-C1 Occupy
We use the pre-trained AnyGrasp[20] to produce a variety of grasp pose proposals.
III-C2 Tool use
We have specifically designed tool-usage skills based on the categories of tools employed, and have provided adjustable parameters to enable the robot to perform flexible tool-manipulating actions.
III-C3 Release
Based on the current grasping position of the object and the 3D bounding box of the target object, we propose an appropriate release pose to ensure that the target object can be accurately placed at the desired location.
III-C4 Operate
We employ the GAPartNet[21] to forecast the physical characteristics of articulated objects. This model works by dividing the point cloud of an articulated object into its constituent rigid parts and subsequently calculating the articulation parameters.
Figure 3: Snapshots of 5 Tasks of Dual-arm Kitchen Benchmark.
TABLE I: Task list of Dual-arm Kitchen Benchmark.
Index Task Name Human Instruction
Task 1 Clean the table (Easy) Clean the table. Put objects into plate.
Task 2 Clean the table (Hard) Clean the table. The fruits should into the plate. The table should be wiped by sponge. The mug and sponge should in drawer.
Task 3 Stack bowls Stack the bowls onto the wooden tray with the green bowl, blue bowl and yellow bowl order.
Task 4 Make cup of coffee Make a cup of coffee. You should add the coffee, water and milk in order. Finally, stir it with the spoon.
Task 5 Boil vegetables Boil vegetables. Pour water and put vegetables into the pot.
IV Experiments
In the experiments, we are interested in answering the following questions about DAG-Plan:
  • $\bullet$
    To what extent can the task-graph approach improve performance compared to the task-sequence method?
  • $\bullet$
    How does the cost of DAG-Plan compare to that of iterative methods?
  • $\bullet$
    How does the performance of DAG-Plan compare to multi-agents methods?
  • $\bullet$
    Can DAG-Plan be deployed in real-world setting?
IV-A Experimental Setup
To validate the correctness and execution efficiency of our method, we created a Dual-arm Kitchen Benchmark, including plan tests and physical simulation tests. This benchmark fills the gap in long-sequence operation physics simulation benchmarks for dual-arm robots. The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning and executable ability in complex scenarios. The dual-arm robot should correctly plan and fully utilize both the left and right arms, completing tasks with as few execution stages as possible. The benchmark consists of 5 sequential tasks are shown in Table I and Figure 3, comprising a total of 44 sub-tasks. Our physical simulation scene is built on the Sapien [34]. The embodied platform is Agilex CobotMagic, each 6 DOF arm equipped with a two-finger gripper. The platform is equipped with an Intel RealSense L515 RGB-D camera, attached on the robot’s head.
IV-A1 Evaluation of Planning Effectiveness and Conciseness
In this experiment, we focus on testing the conciseness of the generated plans and the number of stages required. In a stage, the robot can execute a right arm node and a left arm node. This requires that the plans generated by the LLMs can achieve the task goals in terms of language logic and do not violate the preconditions for stage execution. We used LLMs to generate 10 plans for each task, evaluating their Success Rate (SR) and the minimum Stage of the passed plan required at language level. Due to cost considerations, we also include average Tokens usage as an important metric for evaluating cost of the algorithm. Finally, we calculated the average success rate and the average number of stages for all tasks. We defined Stage Efficiency as the ratio of single-arm plan stages to the stages required by each method. For failed dual-arm plans, we calculated the stage count based on the single-arm plan stages to ensure a fair comparison.
IV-A2 Evaluation with Physical Simulation
In this experiment, we will test the executability and execution efficiency of the plans in physical simulation scenarios. Compared to plan tests, physical simulation tests validate both the high-level planning and low-level execution capabilities. Then evaluate the Success Rate (SR) and minimum Time which consists of query time and execution time. in the physical environment with 10 trials. Finally, we calculated the average success rate and the time for all tasks. We defined Execution Efficiency as the ratio of single-arm plan time to the time required by each method. For failed dual-arm plans, we calculated the average time based on the single-arm time to ensure a fair comparison.
IV-A3 Baseline Algorithms and Proposed Method
We compare the planning algorithms baselines and our proposed method:
  1. 1.
    Task Planning for Single-arm (TP-S): TP-S directly uses LLMs to generate a full task sequence, with each stage involving a single arm to manipulate a single object.
  2. 2.
    Task Planning for Dual-arm (TP-D): TP-D directly uses LLMs to generate a full task sequence, with each stage involving dual arms.
  3. 3.
    Twostep [11]: Twostep decomposes complex tasks into fully independent sub-tasks with dual agents, with the primary agent and the secondary agent each completing separate sub-tasks to accelerate task execution.
  4. 4.
    RoCo [12]: RoCo proposed multi-agents collaboration using LLMs for planning which is an iterative method. Robots query LLMs in each stages to collaboratively reason about task strategies. DABICO [13] is a dual-armed version of the RoCo, and is equivalent to the RoCo in our experiments settings.
  5. 5.
    DAG-Plan (Proposed): Our method DAG-Plan generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
IV-B Evaluation of Planning Effectiveness and Conciseness
Figure 4: Simulation snapshots of the execution process of long-horizon task 2.
TABLE II: Performance comparison on plan tests. We report the success rate, minimum stage and average token usage of the 10 plans generated by the LLM for each task.
Task1 Task2 Task3 Task4 Task5 Average
[…]
TP-S $100\%$ $8$ $2030.0$ $70\%$ $11$ $2424.8$ $100\%$ $6$ $2086.8$ $90\%$ $12$ $2413.0$ $100\%$ $7$ $2096.6$ $92\%$ $100.0\%$ $2210.2$
[…]
TwoStep[11] $100\%$ $4$ $3699.6$ $0\%$ $Fail$ $4248.2$ $0\%$ $Fail$ $3823.4$ $0\%$ $Fail$ $4130.2$ $100\%$ $5$ $3810.2$ $40\%$ $115.8\%$ $3942.3$
RoCo[12] $100\%$ $4$ $26615.0$ $0\%$ $Fail$ $68921.4$ $100\%$ $4$ $25410.2$ $40\%$ $8$ $65242.0$ $100\%$ $4$ $26985.0$ $68\%$ $141.9\%$ $42634.7$
DAG-Plan $100\%$ $4$ $2031.0$ $70\%$ $7$ $2207.2$ $100\%$ $4$ $1980.6$ $70\%$ $7$ $2277.0$ $100\%$ $4$ $1973.0$ $88\%$ $169.2\%$ $2093.8$
TABLE III: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials. Time consists of query time and execution time.
Task1 Task2 Task3 Task4 Task5 Average
[…]
TP-S $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}9.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}53.5}=63.0$ $40\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}12.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}90.7}=103.6$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.6}=43.9$ $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}116.9}=125.1$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.2}+{\color[rgb]{0.78515625,0.11328125,0.19140625}55.8}=64.0$ $78\%$ $100.0\%$
TP-D $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.1}=34.6$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=51.4$ $26\%$ $111.4\%$
TwoStep[11] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.3}=38.7$ $0\%$ $Fail$ $0\%$ $Fail$ $0\%$ $Fail$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}43.2}=54.5$ $26\%$ $109.2\%$
RoCo[12] $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}49.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}26.2}=75.5$ $0\%$ $Fail$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.9}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.1}=79.0$ $30\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}155.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}78.9}=234.2$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}54.4}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.3}=89.7$ $60\%$ $68.7\%$
DAG-Plan $80\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.5}+{\color[rgb]{0.78515625,0.11328125,0.19140625}27.6}=36.1$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}11.7}+{\color[rgb]{0.78515625,0.11328125,0.19140625}59.8}=71.5$ $90\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}24.0}=32.3$ $50\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}69.9}=77.9$ $100\%$ ${\color[rgb]{0.34375,0.55859375,0.19140625}8.3}+{\color[rgb]{0.78515625,0.11328125,0.19140625}35.4}=43.7$ $74\%$ $152.8\%$
As shown in Table II, in the plan tests, DAG-Plan consistently outperformed baselines, showcasing superior efficiency and robustness. DAG-Plan achieved a high success rate across all tasks, demonstrating its effectiveness in dual-arm manipulation. Notably, it maintained an impressive macro average success rate of 88% and exhibited a significant reduction in the required stages for task completion, achieving 169.2% stage efficiency. These results underscore the effectiveness of DAG-Plan in achieving task goals and its efficiency in execution.
In contrast, TP-S, primarily focused on the single-arm plan, generally required more stages to complete tasks compared to TP-D and DAG-Plan. Although TP-S maintained a relatively high and consistent success rate, it was less efficient in stage minimization.
Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 40%), and often produced plans that were not executable in the physical environment. While theoretically capable of reducing the number of stages required for tasks, TP-D frequently encountered challenges related to coordination complexities and unrealistic task assignments. This highlights the superiority of DAG-Plan in effectively translating high-level plans into physical actions and navigating complexities in the environment with relative ease compared to TP-D.
The key idea of Twostep is to decompose a task into two completely independent sub-tasks. While this approach is highly effective in multi-agents task settings, it falls short in dual-arm tasks. In dual-arm scenarios, the entire task sequence is interconnected, especially in tasks with order requirements, such as Task 3 and Task 4. Each sub-task is inherently dependent on others, making it impossible to decompose the task into fully independent sub-tasks. As a result, Twostep achieves low success rates and is ineffective in dual-arm tasks.
RoCo is capable of acquiring the current environment and robot state at each step. Through communication between the left and right arms, it can often select appropriate subtasks, achieving high success rates and phase efficiency for all tasks except Task 2. In Task 2, RoCo is still limited by the drawback of sequentially generating dual-arm plans, neglecting implicit dependency conditions. For instance, it attempts to grasp the sponge and mug without first opening the drawer, resulting in a failure to open the drawer and ultimately causing the plan to fail. Additionally, RoCo requires querying the large model and retaining historical dialogue in each round, leading to significant token consumption in long-sequence tasks.
The success of DAG-Plan can be attributed to its DAG generation, which systematically breaks down tasks into manageable components. This allows for more accurate and feasible plan generation, as the method iteratively refines the task execution steps while maintaining alignment with the operational capabilities of the robotic system. The failed case of DAG-Plan is LLMs occasionally generate some unreasonable task graphs. For instance, each node depend on the previous node, causing the task graph to degenerate into a task sequence, or certain dependencies be omitted.
IV-C Evaluation with Physical Simulation
As shown in Table III, the physical simulation tests provided further insights into the practical applicability and execution capabilities of our planning methods under more dynamic and realistic conditions. Here again, DAG-Plan demonstrated a balanced performance with a solid success rate and efficient times. We show the execution process of DAG-Plan in a physical simulation environment in Figure 4.
Figure 5: Real-world snapshots of the execution process of long-horizon tasks.
Compared to TP-D, DAG-Plan effectively translates high-level plans into feasible actions based on target object information and the robot’s current state under the guidance of a task graph. Both DAG-Plan and TP-S were able to complete $/5$ 5 tasks in the physical simulation tests. However, while TP-D passed the plan tests for $3/5$ tasks, it only completed $2/5$ in the physical simulation tests. In tasks 5, the plans generated by TP-D could not be executed due to real-world physical constraints. In task 5, the plan of TP-D included “grasp red bell pepper” with left arm and “grasp carrot” with right arm. However, since the red bell pepper and carrot both on the table’s right side and too close, the left-arm could not grasp red bell pepper. Instead DAG-Plan select sub-tasks “grasp water bottle” with left arm and “grasp carrot” with right arm. Ultimately, DAG-Plan achieved a 48% higher success rate in physical simulation tests compared to TP-D. This demonstrates that, both in high-level planning and low-level execution, the success rate of the dual-arm plans generated by TP-D is significantly inferior to those generated by DAG-Plan. Furthermore, although RoCo’s stage efficiency is high, it is limited by the fact that each stage needs to be associated with query LLMs, which makes its query time much higher than that of the non-iterative approach, resulting in inefficiencies in actual execution.
Regarding execution efficiency, the dual-arm plan allows for parallel execution of sub-tasks, resulting in higher efficiency. DAG-Plan’s execution efficiency was 52.8% higher than TP-S, as it maximized the parallelization of sub-tasks while ensuring the feasibility of the plan. Despite TP-D had similar execution times to DAG-Plan for some tasks, its overall efficiency was low because $3/5$ tasks could not be executed. Consequently, TP-D’s efficiency was only 11.4% higher than TP-S. Although RoCo has a high success rate and stage efficiency, it is limited by query time which is even longer than execution time, and the execution efficiency is only 68.7%.
IV-D Real-world Experiments
We validate DAG-Plan in real-world setting, where a Agilex Cobot Magic dual-arm robot to complete tasks similar to simulation. When dealing with these tasks, the DAG-Plan framework can effectively assign the sub-task for the each arm. We show the execution process of DAG-Plan in real-world in Figure 5 first row. We evaluate 10 runs for each task as shown in Figure 6 for comparison of real-world and simulation success rates. The slightly lower success rate in real-world is primarily due to the inaccuracy of depth, whereas the depth in simulations is entirely precise.
Figure 6: Comparison of real-world and simulation success rates of DAG-Plan.
In addition to the predefined tasks similar to simulation, we also conducted tests in the wild, as shown in Figure 5 second row. By providing human instructions through audio input, the LLM automatically generates environment descriptions and DAG, enabling the dual-arm robot to efficiently complete the tasks.
V Conclusion
This work introduces DAG-Plan, which efficiently and accurately generates collaborative plans with LLMs for dual-arm robots. DAG-Plan decomposes complex tasks into directed acyclic graph (DAG) with clear temporal relationships and iteratively selects feasible sub-tasks based on environmental observations during execution. The main contribution is the replacement of task sequence with a DAG and the dynamic adjustment of the planning according to the current situation, allowing dual-arm robots to flexibly utilize both arms for sub-task execution. Extensive experiments demonstrate that DAG-Plan integrates the advantages of various methods, exhibiting balanced and superior performance. Compared to the single-arm approach, DAG-Plan demonstrates significantly higher execution efficiency. In contrast to dual-arm methods, DAG-Plan achieves a higher success rate in task planning. When compared to the iterative method, DAG-Plan maintains the ability to interact with environmental information while requiring shorter query time and reduced token usage.
ACKNOWLEDGEMENTS
The authors would like to thank Qiaojun Yu, Tianxing Chen for their fruitful discussions throughout the project and for providing helpful feedback on initial drafts of the manuscript.